OpenStudy (anonymous):

(512^.3)(512^.03)(512^.003) tel me the formula to solve

OpenStudy (zehanz):

512, hmm...this number looks suspicious. Oh it's a power of 2 (512=2^9). Also, this rule of powers is important:\[a^b*a^c=a^{b+c}\]So we get\[512^{.3+.03+.003}=512^{0.333}\] Now, we have a problem: 0.333 is close to 1/3, but it is not the same! Let's assume that 0.333 IS 1/3, then we get:\[512^{\frac{ 1 }{ 3 }}=(2^9)^{\frac{ 1 }{ 3 }}=2^{\frac{ 9 }{ 3 }}=2^3=8\]